Use the following financial statements for Lake of Egypt Marina, Inc.

LAKE OF EGYPT MARINA, INC.
Balance Sheet as of December 31, 2021 and 2020
(in millions of dollars)
2021 2020 2021 2020
Assets Liabilities and Equity
Current assets: Current liabilities:
Cash and marketable securities $ 75 $ 65 Accrued wages and taxes $ 40 $ 43
Accounts receivable 115 110 Accounts payable 90 80
Inventory 200 190 Notes payable 80 70
Total $ 390 $ 365 Total $ 210 $ 193
Fixed assets: Long-term debt: $ 300 $ 280
Gross plant and equipment $ 580 $ 471 Stockholders’ equity:
Less: Depreciation 110 100 Preferred stock (5 million shares) $ 5 $ 5
Net plant and equipment $ 470 $ 371 Common stock and paid-in surplus
(65 million shares) 65 65
Other long-term assets 50 49 Retained earnings 330 242
Total $ 520 $ 420 Total $ 400 $ 312
Total assets $ 910 $ 785 Total liabilities and equity $ 910 $ 785
LAKE OF EGYPT MARINA, INC.
Income Statement for Years Ending December 31, 2021 and 2020
(in millions of dollars)
2021 2020
Net sales (all credit) $ 515 $ 432
Less: Cost of goods sold 230 175
Gross profits $ 285 $ 257
Less: Other operating expenses 30 25
Earnings before interest, taxes, depreciation, and amortization (EBITDA) $ 255 $ 232
Less: Depreciation 22 20
Earnings before interest and taxes (EBIT) $ 233 $ 212
Less: Interest 33 30
Earnings before taxes (EBT) $ 200 $ 182
Less: Taxes 42 55
Net income $ 158 $ 127
Less: Preferred stock dividends $ 5 $ 5
Net income available to common stockholders $ 153 $ 122
Less: Common stock dividends 65 65
Addition to retained earnings $ 88 $ 57
Per (common) share data:
Earnings per share (EPS) $ 2.354 $ 1.877
Dividends per share (DPS) $ 1.000 $ 1.000
Book value per share (BVPS) $ 6.077 $ 4.723
Market value (price) per share (MVPS) $ 14.750 $ 12.550


Calculate the following ratios for Lake of Egypt Marina, Inc. as of year-end 2021. (Use sales when computing the inventory turnover and use total stockholders' equity when computing the equity multiplier. Round your answers to 2 decimal places. Use 365 days a year.)

Profit Margin) (answer is not 30.68)

ROA) (answer is not 17.36 or 18.64)

Days sales in inventory) (the answer is not 309.46 or 138.21)

ROE) (the answer is not 38.25 or 44.38 or 39.50)
Answer #1

Profit Margin= Gross Profit/ Revenue = 285/515 = 55.33%

Return Of Asset = Net Income available to stockholders/ Avg.Total Assets = 153/(910+785)/2 = 18.05%

Days Sale in inventory= 365/ inventory Turnover, Inventory Turnover = COGS/avg Inventory = 230/(190+200)/2 = 1.179

ROE= Net Income available to stockholders/Shareholder Equity = 153/400= 38.25%